use substitution method to determine whether the system of linear equations has no solution, one solution,or many solutions
x + 2y = 0
y = - 1/2x

Answer by robertb(5830)

x + 2y = 0
y = - 1/2x
==> x + 2*(-x/2) = 0 ==> x - x = 0, or 0 = 0, which is a true statement.
There are infinitely-many solutions.